Given are five observations for two variables, x and y. xi 1 2 3 4 5 У|4751216

c. Develop the estimated regression equation by computing the the slope and the y intercept of the estimated regression line (to 1 decimal) d. Use the estimated regression equation to predict the value of y when x- 4 (to 1 decimal)
Answer #1

c
Step 1: Find X.Y and X2 as it was done in the table below 14 15 48 80 4 12 16 16 25 Step 2: Find the sum of every column: Ste
The graph of the regression line is: y=0.1+2.9 x 20 T 18 14 12 10 2

d)
y = 0.1 + 2.9x
for x = 4
y = 0.1 + 2.9(4) = 11.7